# MMTU: A Massive Multi-Task Table Understanding and Reasoning Benchmark

### v1

@@ 1 Introduction
Remarkable progress has been made in foundation models [39, 40, 68, 112, 33], partly thanks to an expanding array of large-scale benchmarking efforts. Prominent examples include benchmarks for general language understanding (e.g., GLUE [114], Super-GLUE [115], BIG-Bench [108], MMLU [73], MMLU-pro [118]), as well as benchmarks focused on coding and STEM reasoning (e.g., SWE-bench [13], GPQA-diamond [103], AIME [5], LiveCodeBench [11]). These efforts have played a critical role in deepening our understanding and accelerating the progress of foundation models.
Tables and table-based use cases are central to many real-world applications, including spreadsheets [21, 2], databases [17, 18], and computational notebooks [25, 19], which often require expert-level users such as data engineers, analysts, and database administrators to operate. While LLMs have shown great promise in working with tables [22, 16, 24], existing evaluations of table-tasks remain narrow in scope – primarily focusing on NL-to-SQL [86, 137, 128] and Table-QA [127, 54, 49] – and fail to reflect the broader spectrum of real-world tasks that professional users face.
@@
Our evaluation shows that the complex and technical nature of MMTU demands a combination of capabilities – including table understanding, reasoning, and coding – that remain challenging for today’s frontier models. Even the top-performing models, such as OpenAI o4-mini and DeepSeek , achieve only 63.9% and 59.6% on MMTU, suggesting substantial room for improvement, and highlighting MMTU as a strong testbed for models aspiring toward general human-level intelligence, e.g., to meet or surpass the top 10% of skilled adults in diverse technical tasks like characterized in [96].
@@
Table Transform table-transform-by-relationalization (TTBR) Relationalize a table using transformation Acc [87, 35, 78] 238
@@
table-transform-by-output-schema (TTBS) Synthesize transformation by output schema Acc [125] 700
@@
table-transform-by-output-table (TTBT) Synthesize transformation by input/output tables Acc [36, 116, 113, 132] 87
@@
Table Matching Entity matching (EM) Match rows refer to the same semantic entity Acc [61, 97, 136, 89, 99] 4380
@@
Schema matching (SM) Match columns refer to the same concept F1 [81, 135, 37, 94] 723
@@
Head value matching (HVM) Match column-headers with cell-values Acc [88, 102] 952
@@
Data Cleaning data-imputation (DI) Predict missing values in tables Acc [41, 64, 88] 2000
@@
error-detection (ED) Detect erroneous cells in tables F1 [56, 117, 95, 46] 1987
@@
list-to-table (L2T) Split lists of undelimited values into table Acc [55, 42, 65, 84] 1000
@@
Table Join semantic-join (SJ) Predict semantic join between two tables Acc [71, 31, 60, 119] 131
@@
equi-join-detect (EJ) Predict equi-joins between a set of tables F1 [90, 52, 105, 76] 517
@@
Column Transform program-transform-by-example (PTBE) Program transformation by input/output examples Acc [72, 66, 63, 107] 568
formula-by-context (FBC) Predict formula based on table context Acc [48, 50] 3626
@@
semantic-transform-by-example (STBE) Predict semantic transformations by examples Acc [71, 31, 60] 131
Column Relationship arithmetic-relationship (AR) Predict arithmetic-relationship (AR) in tables F1 [106, 1] 819
functional-relationship (FR) Predict functional-relationship (FR) in tables F1 [1, 100, 120] 309
string-relationship (SR) Predict string-relationship (SR) in tables F1 [1, 72, 67] 766
@@
Table understanding Needle-in-a-haystack-table (NIHT) Retrieve cell content in a table Acc [64] 1000
@@
Needle-in-a-haystack-index (NIHI) Retrieve index based on cell value Acc [64] 1000
NL-2-code NL-2-SQL (NS) Translate natural-language into SQL Acc [137, 86, 83, 128, 138] 3289
@@
Table QA Table Question Answering (TQA) Answer questions based on tables Acc [122, 127, 53, 54] 2424
@@
Fact Verification (FV) Verify facts based on tables Acc [49, 126, 131] 1000
@@
KB Mapping Column type annotation (CTA) Predict KB types based on column content Acc [77, 124, 109, 130, 75] 1000
@@
Column property annotation (CPA) Predict KB property for a pair of columns Acc [77, 59, 98] 1000
@@
Cell entity annotation (CEA) Predict KB entity for a table cell Acc [77, 62, 38, 121, 92] 1000
@@ 2 Related Work
Large-scale benchmarks for foundation models. The rapid advancement of foundation models has made benchmark evaluation ever more important. Prominent large-scale benchmarks, such as GLUE [114] (9 NLP tasks), Super-GLUE [115] (10 NLP tasks), Big-BENCH [108] (204 tasks), MMMU [73] (15,908 questions), MMMU-pro [118] (12,032 questions), MMLU [129] (11,550 multi-modal questions) etc., offer comprehensive evaluations of model capabilities. However, as models improve rapidly, benchmarks can become saturated quickly (e.g., in the matter of a few years), prompting newer and more challenging benchmarks [115, 118]. All of these benchmarking efforts have nevertheless played a crucial role in measuring and stimulating the development of foundation models.
@@
Benchmarks for reasoning. Reasoning has recently emerged as a key challenge for foundation models. Beyond general intelligence benchmarks (e.g., GPQA Diamond [103], AGIEval [139], HLE [8]), there are specialized benchmarks targeting mathematical reasoning (e.g., AIME [5], MathVista [12], IMO [9]), coding (e.g., SWE-bench [13], CodeForce [7], LiveCodeBench [11], IOI [10]), and multi-modal reasoning (e.g., MMMU [129], ARC [6]). These benchmarks have become important tools for evaluating models’ ability to tackle complex reasoning tasks, but can also get saturated quickly (e.g., GSM8k [57], Math500 [74], HuamEval [45]), making it necessary to create new and more challenging benchmarks.
@@
Existing benchmarks relating to tables. Given the importance of table data, benchmarks have been developed in the ML and NLP community to evaluate model ability on tables, which however usually focus on a small set of table tasks such as NL-2-SQL [137, 86, 83, 128, 138] and Table-QA [49, 122, 127, 53, 54]. In contrast, MMTU expands the scope of current evaluations by incorporating 19 additional table tasks drawn from decades of research in communities such as data management and programming languages, resulting in a more comprehensive evaluation framework for assessing LLM capabilities on tabular data.
More recently, spreadsheet-centric benchmarks have emerged, including Spreadsheet-Bench [93], SheetCopilotBench [85], and Sheet-RM [51]. While these are important in the spreadsheet domain, these efforts are typically limited in scale (containing a few hundred cases), and are closely tied to specific file formats (e.g., .xlsx) and software environments. In comparison, MMTU focuses on general-purpose tabular data that applies broadly across spreadsheet, database, and computational notebook settings, enabling more scalable and format-independent evaluation of foundation models.
@@ 3.2 Data curation workflow
Literature survey. To ensure our table tasks reflect real-world challenges, we draw on our experience working on related problems, that many challenging predictive table tasks have been studied in the decades worth of computer science research, particularly in data management (SIGMOD/VLDB), programming languages (PLDI/POPL), and web data (WWW/WSDM) communities. We conduct a systematic survey of publications from these venues over the past two decades, leveraging a combination of keyword searches of paper titles, and DeepResearch-like tools [20] (where we specify detailed requirements for papers in these venues), to arrive at a promising set of papers and possible candidate tasks.
@@
  • (1) Real user-facing tasks, involving data tables that would otherwise require expert-level humans to perform. (We therefore exclude system-level predictive table tasks focused on performance improvements, such as query optimization [80, 82] and cardinality estimation [69, 79]);
  • (2) Objectively evaluable tasks, that come with unique manually-labeled ground truth. (We therefore exclude tasks such as table summarization [134, 34, 70, 101, 44] and table augmentation [133, 123], which lack unique ground-truth and may require subjective fuzzy LLM-based evaluations);
@@
Evaluation framework. In contrast to benchmarks like MMLU [73] and MMMU [129], which primarily use multiple-choice formats (where evaluation involves comparing a predicted option such as “A/B/C/D” against a single-letter ground truth), real-world table tasks performed by professional experts are often far more complex and nuanced. These tasks, such as code generation or structured reasoning, cannot be adequately evaluated using multiple-choice alone. In MMTU, we instead adopt a structured yet open-ended answer format (see Figure 1 for examples) for prediction and evaluation.
To support evaluations beyond simple string comparisons, we designed a lightweight evaluation framework that supports diverse evaluations in table tasks, including execution-based evaluation (for SQL and Python generation), and structured output evaluation (e.g., comparing an unordered JSON list against ground truth). Our framework is also extensible, making it easy to incorporate new task types and evaluation metrics. Details of our evaluation framework can be seen in [15].
@@ 4 Experiments
Experimental setup. In all our experiments reported below, we use publicly available model endpoints for inference [4, 23] with default parameter settings. All of our code and data are publicly available at [15] for future research.
@@ 4.1 Overall performance
Chat GPT-4o (2024-11-20) 0.491 ± 0.01 0.0204
GPT-4o-mini (2024-07-18) 0.386 ± 0.01 0.0024
Llama-3.3-70B 0.438 ± 0.01 0.0027
Llama-3.1-8B 0.259 ± 0.01 0.0008
Mistral-Large-2411 0.430 ± 0.01 0.0108
Mistral-Small-2503 0.402 ± 0.01 0.0006
Reasoning o4-mini (2024-07-18) 0.639 ± 0.01 0.0099
Deepseek-R1 0.596 ± 0.01 0.0052
Table 3: Overall performance and cost results of frontier models.
@@
From a cost-efficiency perspective [^1] Price-per-token figures are sourced from the respective API providers [27, 28, price-gpt-4o-mini, price-llama3.1-8B, price-llama3.3-70B, price-mistral-large-2411, price-mistral-small-2503, 30]., both reasoning models are also competitive (despite producing more intermediate thinking tokens), due to the relatively light-weight nature of R1 and o4-mini, making them cost-effective choices for tackling table tasks.

### v2

@@ 1 Introduction
Remarkable progress has been made in foundation models Bommasani et al. (2021); Brown et al. (2020); Guo et al. (2025); Touvron et al. (2023); Anil et al. (2023), partly thanks to an expanding array of large-scale benchmarking efforts. Prominent examples include benchmarks for general language understanding (e.g., GLUE Wang et al. (2018), Super-GLUE Wang et al. (2019), BIG-Bench Srivastava et al. (2022), MMLU Hendrycks et al. (2020), MMLU-pro Wang et al. (2024)), as well as benchmarks focused on coding and STEM reasoning (e.g., SWE-bench ben (i), GPQA-diamond Rein et al. (2024), AIME ben (a), LiveCodeBench ben (g)). These efforts have played a critical role in deepening our understanding and accelerating the progress of foundation models.
Tables and table-based use cases are central to many real-world applications, including spreadsheets exc (a); Goo , databases dat (b, c), and computational notebooks jup ; dat (d), which often require expert-level users such as data engineers, analysts, and database administrators to operate. While LLMs have shown great promise in working with tables exc (b); dat (a); gem , existing evaluations of table-tasks remain narrow in scope – primarily focusing on NL-to-SQL Li et al. (2024); Zheng et al. (2024); Yu et al. (2018) and Table-QA Yang et al. (2015); Cheng et al. (2021); Chen et al. (2019) – and fail to reflect the broader spectrum of real-world tasks that professional users face.
@@
Our evaluation shows that the complex and technical nature of MMTU demands a combination of capabilities – including table understanding, reasoning, and coding – that remain challenging for today’s frontier models. Even the top-performing models, such as OpenAI o4-mini and DeepSeek , achieve only 63.9% and 59.6% on MMTU, suggesting substantial room for improvement, and highlighting MMTU as a strong testbed for models aspiring toward general human-level intelligence, e.g., to meet or surpass the top 10% of skilled adults in diverse technical tasks like characterized in Morris et al. (2023).
@@
Table Transform table-transform-by-relationalization (TTBR) Relationalize a table using transformation Acc Li et al. (2023b); Barowy et al. (2015); Jin et al. (2017) 238
@@
table-transform-by-output-schema (TTBS) Synthesize transformation by output schema Acc Yang et al. (2021) 700
@@
table-transform-by-output-table (TTBT) Synthesize transformation by input/output tables Acc Bavishi et al. (2019); Wang et al. (2017); Tran et al. (2009); Zhang and Sun (2013) 87
@@
Table Matching Entity matching (EM) Match rows refer to the same semantic entity Acc Das et al. ; Mudgal et al. (2018); Zhao and He (2019); Li et al. (2020); Papadakis et al. (2021) 4380
@@
Schema matching (SM) Match columns refer to the same concept F1 Koutras et al. (2021); Zhang et al. (2024); Bellahsene et al. (2011); Madhavan et al. (2001) 723
@@
Head value matching (HVM) Match column-headers with cell-values Acc Li et al. (2023c); Rahm and Bernstein (2001) 952
@@
Data Cleaning data-imputation (DI) Predict missing values in tables Acc Cafarella et al. (2008a); Eggert et al. (2023); Li et al. (2023c) 2000
@@
error-detection (ED) Detect erroneous cells in tables F1 Chu et al. (2016); Wang and He (2019); Mahdavi and Abedjan (2020); Chen et al. (2025) 1987
@@
list-to-table (L2T) Split lists of undelimited values into table Acc Chu et al. (2015); Cafarella et al. (2008b); Elmeleegy et al. (2009); Lerman et al. (2001) 1000
@@
Table Join semantic-join (SJ) Predict semantic join between two tables Acc He et al. (2015); Abedjan et al. (2016); Dargahi Nobari and Rafiei (2024); Wang and He (2017) 131
@@
equi-join-detect (EJ) Predict equi-joins between a set of tables F1 Lin et al. (2023); Chen et al. (2014); Rostin et al. (2009); Jiang and Naumann (2020) 517
@@
Column Transform program-transform-by-example (PTBE) Program transformation by input/output examples Acc He et al. (2018); Gulwani (2011); Devlin et al. (2017); Singh (2016) 568
formula-by-context (FBC) Predict formula based on table context Acc Chen et al. (2024a, 2021b) 3626
@@
semantic-transform-by-example (STBE) Predict semantic transformations by examples Acc He et al. (2015); Abedjan et al. (2016); Dargahi Nobari and Rafiei (2024) 131
Column Relationship arithmetic-relationship (AR) Predict arithmetic-relationship (AR) in tables F1 Shraga and Miller (2023); Aut 819
functional-relationship (FR) Predict functional-relationship (FR) in tables F1 Aut ; Parciak et al. (2024); Wei and Link (2019) 309
string-relationship (SR) Predict string-relationship (SR) in tables F1 Aut ; He et al. (2018); Gulwani et al. (2012) 766
@@
Table understanding Needle-in-a-haystack-table (NIHT) Retrieve cell content in a table Acc Eggert et al. (2023) 1000
@@
Needle-in-a-haystack-index (NIHI) Retrieve index based on cell value Acc Eggert et al. (2023) 1000
NL-2-code NL-2-SQL (NS) Translate natural-language into SQL Acc Zheng et al. (2024); Li et al. (2024); Lee et al. (2021); Yu et al. (2018); Zhong et al. (2017) 3289
@@
Table QA Table Question Answering (TQA) Answer questions based on tables Acc Wu et al. (2025); Yang et al. (2015); Chen et al. (2021c); Cheng et al. (2021) 2424
@@
Fact Verification (FV) Verify facts based on tables Acc Chen et al. (2019); Yang and Zhu (2021); Zhang et al. (2020a) 1000
@@
KB Mapping Column type annotation (CTA) Predict KB types based on column content Acc Jiménez-Ruiz et al. (2020); Yan and He (2018); Suhara et al. (2022); Zhang et al. (2019); Hulsebos et al. (2019) 1000
@@
Column property annotation (CPA) Predict KB property for a pair of columns Acc Jiménez-Ruiz et al. (2020); Cutrona et al. (2022); Nguyen et al. (2021) 1000
@@
Cell entity annotation (CEA) Predict KB entity for a table cell Acc Jiménez-Ruiz et al. (2020); Deng et al. (2022); Bhagavatula et al. (2015); Wu et al. (2016); Liu et al. (2023) 1000
@@ 2 Related Work
Large-scale benchmarks for foundation models. The rapid advancement of foundation models has made benchmark evaluation ever more important. Prominent large-scale benchmarks, such as GLUE Wang et al. (2018) (9 NLP tasks), Super-GLUE Wang et al. (2019) (10 NLP tasks), Big-BENCH Srivastava et al. (2022) (204 tasks), MMMU Hendrycks et al. (2020) (15,908 questions), MMMU-pro Wang et al. (2024) (12,032 questions), MMLU Yue et al. (2024) (11,550 multi-modal questions) etc., offer comprehensive evaluations of model capabilities. However, as models improve rapidly, benchmarks can become saturated quickly (e.g., in the matter of a few years), prompting newer and more challenging benchmarks Wang et al. (2019, 2024). All of these benchmarking efforts have nevertheless played a crucial role in measuring and stimulating the development of foundation models.
@@
Benchmarks for reasoning. Reasoning has recently emerged as a key challenge for foundation models. Beyond general intelligence benchmarks (e.g., GPQA Diamond Rein et al. (2024), AGIEval Zhong et al. (2023), HLE ben (d)), there are specialized benchmarks targeting mathematical reasoning (e.g., AIME ben (a), MathVista ben (h), IMO ben (e)), coding (e.g., SWE-bench ben (i), CodeForce ben (c), LiveCodeBench ben (g), IOI ben (f)), and multi-modal reasoning (e.g., MMMU Yue et al. (2024), ARC ben (b)). These benchmarks have become important tools for evaluating models’ ability to tackle complex reasoning tasks, but can also get saturated quickly (e.g., GSM8k Cobbe et al. (2021), Math500 Hendrycks et al. (2021), HuamEval Chen et al. (2021a)), making it necessary to create new and more challenging benchmarks.
@@
Existing benchmarks relating to tables. Given the importance of table data, benchmarks have been developed in the ML and NLP community to evaluate model ability on tables, which however usually focus on a small set of table tasks such as NL-2-SQL Zheng et al. (2024); Li et al. (2024); Lee et al. (2021); Yu et al. (2018); Zhong et al. (2017) and Table-QA Chen et al. (2019); Wu et al. (2025); Yang et al. (2015); Chen et al. (2021c); Cheng et al. (2021). In contrast, MMTU expands the scope of current evaluations by incorporating 19 additional table tasks drawn from decades of research in communities such as data management and programming languages, resulting in a more comprehensive evaluation framework for assessing LLM capabilities on tabular data.
More recently, spreadsheet-centric benchmarks have emerged, including Spreadsheet-Bench Ma et al. (2024), SheetCopilotBench Li et al. (2023a), and Sheet-RM Chen et al. (2024b). While these are important in the spreadsheet domain, these efforts are typically limited in scale (containing a few hundred cases), and are closely tied to specific file formats (e.g., .xlsx) and software environments. In comparison, MMTU focuses on general-purpose tabular data that applies broadly across spreadsheet, database, and computational notebook settings, enabling more scalable and format-independent evaluation of foundation models.
@@ 3.2 Data curation workflow
Literature survey. To ensure our table tasks reflect real-world challenges, we draw on our experience working on related problems, that many challenging predictive table tasks have been studied in the decades worth of computer science research, particularly in data management (SIGMOD/VLDB), programming languages (PLDI/POPL), and web data (WWW/WSDM) communities. We conduct a systematic survey of publications from these venues over the past two decades, leveraging a combination of keyword searches of paper titles, and DeepResearch-like tools dee (where we specify detailed requirements for papers in these venues), to arrive at a promising set of papers and possible candidate tasks.
@@
  • (1) Real user-facing tasks, involving data tables that would otherwise require expert-level humans to perform. (We therefore exclude system-level predictive table tasks focused on performance improvements, such as query optimization Kossmann et al. (2022); Lan et al. (2021) and cardinality estimation Han et al. (2021); Kim et al. (2022));
  • (2) Objectively evaluable tasks, that come with unique manually-labeled ground truth. (We therefore exclude tasks such as table summarization Zhang et al. (2020b); Bao et al. (2018); Hancock et al. (2019); Parikh et al. (2020); Chen et al. (2013) and table augmentation Zhang and Balog (2017); Yakout et al. (2012), which lack unique ground-truth and may require subjective fuzzy LLM-based evaluations);
@@
Evaluation framework. In contrast to benchmarks like MMLU Hendrycks et al. (2020) and MMMU Yue et al. (2024), which primarily use multiple-choice formats (where evaluation involves comparing a predicted option such as “A/B/C/D” against a single-letter ground truth), real-world table tasks performed by professional experts are often far more complex and nuanced. These tasks, such as code generation or structured reasoning, cannot be adequately evaluated using multiple-choice alone. In MMTU, we instead adopt a structured yet open-ended answer format (see Figure 1 for examples) for prediction and evaluation.
To support evaluations beyond simple string comparisons, we designed a lightweight evaluation framework that supports diverse evaluations in table tasks, including execution-based evaluation (for SQL and Python generation), and structured output evaluation (e.g., comparing an unordered JSON list against ground truth). Our framework is also extensible, making it easy to incorporate new task types and evaluation metrics. Details of our evaluation framework can be seen in cod .
@@ 4 Experiments
Experimental setup. In all our experiments reported below, we use publicly available model endpoints for inference aoa ; fou with default parameter settings. All of our code and data are publicly available at cod for future research.
@@ 4.1 Overall performance
Chat Qwen3-235B-A22B-Instruct-2507 0.509 ± 0.01 0.00044
GPT-4o (2024-11-20) 0.491 ± 0.01 0.01019
Llama-4-Maverick-17B-128E-Instruct-FP8 0.472 ± 0.01 0.00066
Llama-3.3-70B 0.438 ± 0.01 0.0015
Mistral-Large-2411 0.430 ± 0.01 0.01066
Mistral-Small-2503 0.402 ± 0.01 0.00273
GPT-4o-mini (2024-07-18) 0.386 ± 0.01 0.00061
Llama-4-Scout-17B-16E-Instruct 0.377 ± 0.01 0.00036
Qwen3-32B (no thinking) 0.367 ± 0.01 0.00007
Qwen3-8B (no thinking) 0.343 ± 0.01 0.00015
Qwen2.5-7B-Instruct 0.300 ± 0.01 0.00010
Llama-3.1-8B 0.259 ± 0.01 0.00003
Reasoning o3 0.671 ± 0.01 0.01539
Gemini-2.5-pro 0.646 ± 0.01 0.00790
o4-mini (2024-07-18) 0.639 ± 0.01 0.00993
Grok-3-mini 0.624 ± 0.01 0.00111
Gemini 2.5 Flash 0.606 ± 0.01 0.00199
Deepseek-R1 0.596 ± 0.01 0.00167
Grok-3 0.576 ± 0.01 0.01232
gpt-oss-120b 0.523 ± 0.01 0.00050
Qwen3-235B-A22B-Thinking-2507 0.510 ± 0.01 0.00068
Qwen3-32B (thinking) 0.488 ± 0.01 0.00017
gpt-oss-20b 0.460 ± 0.01 0.00040
Qwen3-8B (thinking) 0.460 ± 0.01 0.00041
Table 3: Overall performance and cost results of chat and reasoning models. Cost per question are calculated based on public pricing information on https://openrouter.ai/ as of Aug 18, 2025.
@@
From a cost-efficiency perspective [^1] Price-per-token figures are sourced from the respective API providers pri (a, b, c, d, e, f, g, h)., both reasoning models are also competitive (despite producing more intermediate thinking tokens), due to the relatively light-weight nature of R1 and o4-mini, making them cost-effective choices for tackling table tasks.
@@
Task 0–25% range 25–50% range 50–75% range 75–100% range
Arithmetic-Relationship 1150.0–3547.0 3547.0–6834.0 6834.0–15601.5 15601.5–126654.0
Cell-entity-annotation 263.0–2138.5 2138.5–3706.0 3706.0–4693.5 4693.5–7526.0
Column-type-annotation 274.0–576.8 576.8–1346.5 1346.5–3094.8 3094.8–6744.0
Columns-property-anotation 261.0–612.8 612.8–1416.0 1416.0–3004.2 3004.2–6754.0
Data-Imputation 214.0–628.2 628.2–1320.0 1320.0–1971.0 1971.0–102929.0
Data-transform-pbe 262.0–309.8 309.8–333.5 333.5–412.2 412.2–21846.0
Data-transform-reshape 534.0–916.5 916.5–1692.0 1692.0–3621.0 3621.0–76096.0
Entity-Matching 175.0–207.0 207.0–232.0 232.0–248.5 248.5–425.0
Error-Detect 211.0–235.0 235.0–293.0 293.0–638.5 638.5–21040.0
Formula-prediction-context 269.0–1096.2 1096.2–1315.0 1315.0–1564.0 1564.0–10332.0
Functional-Dependency 1341.0–3641.0 3641.0–5157.0 5157.0–7573.0 7573.0–76083.0
List-to-table 184.0–283.5 283.5–352.0 352.0–514.5 514.5–2955.0
NL2SQL 418.0–872.0 872.0–1245.0 1245.0–3293.0 3293.0–35250.0
Schema-Matching 547.0–1803.5 1803.5–2378.0 2378.0–3737.5 3737.5–10562.0
String-Relationship 2961.0–7596.2 7596.2–14127.5 14127.5–26032.5 26032.5–92595.0
Table-Fact-Verification 269.0–451.0 451.0–584.0 584.0–836.5 836.5–2959.0
Table-Locate-by-Row-Col 11190.0–21158.0 21158.0–22433.0 22433.0–23731.0 23731.0–72074.0
Table-QA 229.0–381.0 381.0–537.0 537.0–847.2 847.2–13752.0
Table-needle-in-a-haystack 11123.0–21286.0 21286.0–22737.0 22737.0–23664.0 23664.0–72009.0
Transform-by-input-output-table 200.0–259.0 259.0–297.0 297.0–366.0 366.0–728.0
Transform-by-output-target-schema 360.0–1095.5 1095.5–1827.5 1827.5–2845.8 2845.8–113819.0
equi-join-detect 477.0–2471.0 2471.0–3972.0 3972.0–6741.0 6741.0–107690.0
header-value-matching 186.0–260.0 260.0–338.5 338.5–489.2 489.2–1102.0
semantic-join 243.0–367.5 367.5–742.0 742.0–1860.5 1860.5–19184.0
semantic-transform 209.0–247.0 247.0–265.0 265.0–285.0 285.0–454.0
Table 4: Exact ranges of token count calculated for MMTU questions, in each quartile and for each task, as used in long-table-context analysis (Figure 6). The number of tokens is calculated using GPT-4o tokenizer.